Find the coordinates of the point where the line with parametric equations:
x=1-t, y=3+t and z=3-2t meets the X0Y plane.
Thanks :)

Answer by Edwin McCravy(20054)

Find the coordinates of the point where the 
line with parametric equations: 

x=1-t, y=3+t and z=3-2t meets the X0Y plane.

Thanks :)

This is when z=0

x=1-t, y=3+t and z=3-2t

Substitute z = 0

                 z=3-2t
                 0=3-2t
                2t=3
                 t=3/2

Substitute t=3/2

x=1-t,     y=3+t      and z=3-2t
x=1-3/2,   y=3+3/2,   and z=3-2(3/2)
x=2/2-3/2, y=6/2+3/2, and z=3-3
x=-1/2,    y=9/2,     and z=0

So the point where the
line x=1-t, y=3+t and z=3-2t
intersects the xy-plane is

(-1/2, 9/2, 0)

Edwin